3. Given the normally distributed variable X with mean 18 and standard deviation 2.5, find

(a) P(X<15):

(h) P(17< X<21);

(c) the value of k such that P(X<k) = 0.2578;

(d) the value of k such that P(X> A) = 0.1539.


1
Expert's answer
2021-07-05T15:03:05-0400

"\mu= 18 \\

\sigma= 2.5"

(a)

"P(X<15) = P(Z< \frac{15-18}{2.5}) \\

= P(Z< -1.2) \\

= 0.1151"

(b)

"P(17<X<21) = P(X<21) -P(X<17) \\

=P(Z< \frac{21-18}{2.5}) -P(Z< \frac{17-18}{2.5}) \\

= P(Z<1.2) -P(Z< -0.4) \\

= 0.8849 -0.3446 \\

= 0.5403"

(c)

"P(X<k) = 0.2578 \\

P(Z< -0.65) = 0.2578 \\

\frac{k-18}{2.5} = -0.65 \\

k = 18 - 1.65 \times 2.5 \\

k = 13.875"

(d)

"P(X>k) = 0.1539 \\

1 -P(X<k) = 0.1539 \\

P(X<k) = 1 -0.1539 \\

P(X<k) = 0.8461 \\

P(Z<1.02) = 0.8461 \\

\frac{k-18}{2.5}= 1.02 \\

k= 18 + 1.02 \times 2.5 \\

k = 20.55"